# An information-matching approach to optimal experimental design and active learning

### v4

@@ Abstract
We use the IEEE 39-bus system [18], represented graphically in Fig. 2, as a benchmark. Nodes denote buses, and edges represent transmission lines and transformers [53]. The parameters $\boldsymbol{\theta}$ to be inferred are the system states (voltage magnitude and angle), and $\bm{f}$ represents PMU measurements. Full parameter identifiability implies that the QoIs are the state variables themselves $\bm{g}(\boldsymbol{\theta};\mathbf{y})=\boldsymbol{\theta}$] and the QoI FIM is non-singular. We enforce this condition by setting $\mathcal{J}=\lambda I$ for some small $\lambda>0$. The information-matching condition implies that $\mathcal{I}$ is also non-singular, and all states are observable. The optimal PMU placements for this problem are represented as the orange highlighted buses in Fig. 2. Our results on this initial benchmark test align with previous studies [30, 2, 53]. Even without preassigning PMU locations, this method naturally selects the same buses identified elsewhere [28, 11].
@@
Next, we consider an optimal sensor placement problem in passive acoustic source localization in the ocean. The objective is to determine the optimal sensor (sound receiver) locations to infer the location of sound sources in a shallow ocean. This problem is difficult because sound propagation depends in complicated ways on unknown properties of the ocean environment, including the water temperature and the sediments in the seabed. Traditionally, an estimate of the environmental parameters of the ocean is needed to localize the source, as in matched-field processing [41]. In contrast, information-matching allows us to learn only those combinations of acoustic parameters that are necessary to infer the source location.
@@
We aim to localize two vertically separated sound sources at depths of 8 and 16 m (red speakers) within $\pm$ 2.5 m vertically and $\pm$ 100 m horizontally (Fig. 3). Candidate receivers are arranged in a rectangular grid (small dots), motivated by common practices of using vertical and horizontal line arrays for ocean sound measurements [52, 16, 6, 15, 8]. We use a range-independent normal mode model, called ORCA [51], to simulate the sound propagation in the ocean and compute the transmission loss at 200 Hz for each candidate location illustrated as small dots in Fig. 3 (details in the SM). The model parameters $\boldsymbol{\theta}$ include the source and receiver locations, as well as the parameterization of the ocean environment. The ocean environment is modeled with 75 m-deep water above an ocean floor consisting of a sandy sediment layer on a half-space basement layer.
To localize the two sources using Eq. (4), the QoI FIM $\mathcal{J}$ is a diagonal matrix, where diagonal elements that correspond to source locations are set to their inverse target precision, and all other elements are zero. The optimal receiver locations are shown as the large dots in Fig. 3, which account for only 5% of the candidate locations.
@@
In summary, we have introduced an information-matching method to identify a minimal set of informative data points to meet target precision requirements. This approach involves solving the convex optimization problem in Eq. (4) to align the information available in the training data with that necessary for precise predictions. Unlike other FIM-based OED criteria, this method selectively constrains only the parameters relevant for the downstream QoIs, making it robust to numerical noise in sloppy models with unidentifiable parameters and small FIM eigenvalues. The approach not only improves data efficiency but also enhances model interpretability.
We have demonstrated the versatility of the approach across diverse domains, including power system networks, underwater acoustics, and AL for interatomic potential development. We have also explored the hypothesis that relatively few data points may efficiently constrain typical QoIs. The potential applications of this method extend beyond these areas, encompassing domains such as biology, neuroscience, geology, and atmospheric science.

### v5

@@ Abstract
We use the IEEE 39-bus system [18], represented graphically in Fig. 2, as a benchmark. Nodes denote buses, and edges represent transmission lines and transformers [53]. The parameters $\boldsymbol{\theta}$ to be inferred are the system states (voltage magnitude and angle), and $\bm{f}$ represents PMU measurements. Full parameter identifiability implies that the QoIs are the state variables themselves $\bm{g}(\boldsymbol{\theta};\mathbf{y})=\boldsymbol{\theta}$] and the QoI FIM is non-singular. We enforce this condition by setting $\mathcal{J}=\lambda I$ for some small $\lambda>0$. The information-matching condition implies that $\mathcal{I}$ is also non-singular, and all states are observable. The optimal PMU placements for this problem are represented as the orange highlighted buses in Fig. 2. On this initial benchmark test, the information-matching method naturally selects the same buses identified in previous studies [28, 11], even without preassigning PMU locations. Furthermore, this result is broadly consistent with prior findings [30, 2, 53].
@@
Next, we consider an optimal sensor placement problem in passive acoustic source localization in the ocean. The objective is to determine the optimal sensor (sound receiver) locations to infer the location of sound sources in a shallow ocean. This problem is difficult because sound propagation depends in complicated ways on unknown properties of the ocean environment, including the water temperature and the sediments in the seabed. Traditionally, source localization requires first estimating the ocean environmental parameters, as is done in matched-field processing [41]. In contrast, information-matching allows us to learn only those combinations of acoustic parameters that are necessary to infer the source location, avoiding full environmental inversion when it is not required.
We aim to localize two vertically separated sound sources at depths of 8 and 16 m (red speakers) within $\pm$ 2.5 m vertically and $\pm$ 100 m horizontally (Fig. 3). Candidate receivers are arranged in a rectangular grid (small dots), motivated by common practices of using vertical and horizontal line arrays for ocean sound measurements [52, 16, 6, 15, 8]. We use a range-independent normal mode model, called ORCA [51], to simulate the sound propagation in the ocean and compute the transmission loss at 200 Hz for each candidate location illustrated as small dots in Fig. 3 (details in the SM). The model parameters $\boldsymbol{\theta}$ for simulating sound propagation include the source and receiver locations, as well as the parameterization of the ocean environment. The ocean environment is modeled with 75 m-deep water above an ocean floor consisting of a sandy sediment layer on a half-space basement layer.
To localize the two sources using Eq. (4), the QoI FIM $\mathcal{J}$ is a diagonal matrix, where diagonal elements that correspond to source locations are set to their inverse target precision, and all other elements are zero, indicating that only the relevant environmental parameters for localizing the sources are constrained as needed. The optimal receiver locations are shown as the large dots in Fig. 3, which account for only 5% of the candidate locations.
@@
In summary, we have introduced an information-matching method to identify a minimal set of informative data points to meet target precision requirements for downstream QoIs. We have demonstrated the versatility of the approach across diverse domains, including power system networks, underwater acoustics, and AL for interatomic potential development. Unlike classical FIM-based OED criteria that aim at reducing global parameter uncertainty, information-matching optimizes a fundamentally different objective: it matches the information content of the data to the precision required for the QoIs. As a result, only those parameters that influence the QoI are constrained—and only to the extent needed to satisfy the target QoI precision—while others may remain unconstrained if they are irrelevant. This distinction highlights that comparisons with traditional OED methods must be interpreted carefully, as the two frameworks prioritize different aspects of the inference problem and therefore tend to yield different designs. The accompanying theorem additionally formalizes this perspective and shows that QoI precision can be controlled without explicitly estimating the full parameter uncertainty. This opens new possibilities in settings where full parameter identifiability is unnecessary or infeasible, such as PMU placement for partially observable networks or localization of ocean acoustic sources without environmental inversion. Broader applications are likely in areas such as biology, neuroscience, geology, and atmospheric science, where models often contain many weakly identifiable parameters but well-defined QoIs [33].